# Universal mechanistic rules for de novo design of enzymes

### v1

@@ Abstract
Enzymes are nano-scale machines that have evolved to drive chemical reactions out of equilibrium in the right place at the right time. Thermodynamically favourable reactions such as ATP hydrolysis are used by the cell to convert chemical energy into useful structure, function, and mechanical work. This includes the ‘fuelled’ catalysis of chemical reactions that would otherwise be thermodynamically unfavourable. Given the complexity and specificity of enzymatic function, bottom-up design of enzymes presents a daunting task that is far more challenging than making passive molecules with specific binding affinities or building nano-scale mechanically active devices. Here, we present a thermodynamically-consistent model for the operation of such a fuelled enzyme, which uses the energy from a favourable reaction to undergo non-equilibrium conformational changes that in turn catalyze a chemical reaction on an attached substrate molecule. We show that enzymatic function can emerge through a bifurcation upon appropriate implementation of momentum conservation on the effective reaction coordinates of the low dimensional description of the enzyme, and thanks to a generically present dissipative coupling. By considering the different aspects of the dynamics, such as the interplay of the non-equilibrium drive and the geometry of the enzyme-substrate complex, we propose three golden rules that should be universally applicable for de novo design of enzymes, as they are based on generic ingredients and physical constraints. These rules lead to optimal combinations of parameters, which can vastly accelerate reactions, while at the same time decreasing the energy dissipation of the combined reaction process, or, in other words, to an efficient enzyme. Our results can complement the recently developed strategies for de novo enzyme design based on machine learning approaches.
@@ I Introduction
In recent years, fundamental insight from statistical physics has been successfully used towards experimental demonstrations of many interesting non-equilibrium phenomena [15, 16, 17, 18]. To aim towards the development of systems that can accomplish relatively complex tasks, one can benefit from innovative ideas that build on emergent physical properties of many-body non-equilibrium systems. Recent examples of such developments include proposals to employ non-reciprocal interactions to design collective barrier-crossing strategies that could emulate enzymatic function at larger scales [19], autonomous multifarious self-organization of complex protein structures [20], as well as proposed scenarios for fast and efficient self-organization of primitive metabolic cycles at the early stages of life formation [21].
@@
Here, we set out to construct a minimal model for a ‘fuelled’ enzyme, so that we can extract a set of golden rules regarding its optimal design (see below). Our theoretical framework is built upon two main pillars, namely, momentum conservation as a fundamental physical constraint on the effective reaction coordinates that span the relevant low dimensional configuration space of the system, and a generically present dissipative coupling between the different reaction coordinates, which we systematically derive from a microscopic model. The implementation of momentum conservation leads to the emergence of an effective mechanochemical coupling in the same spirit as models used in studies of stochastic nano-swimmers [29, 30, 31, 32] and enhanced diffusion of enzymes [33, 34], while the dissipative coupling has been recently used to show how enzyme pairs can cooperate by exhibiting synchronization and enhanced catalytic activity [35]. The two ingredients described above give rise to a bifurcation in the dynamical phase-portrait of the low dimensional configuration space, which enables enzymatic activity as an emergent feature of the dynamics, via a fundamentally novel mechanism not accessible to Kramers-like energy barrier-crossing descriptions in terms of reaction coordinates. Our work complements other studies that involve minimal models for studying enzymatic behaviour, often based on colloids interacting through short-ranged interaction potentials or bead-spring networks [36, 37, 38, 39, 40, 41], which have been shown to be able to describe the action of a ‘passive’ enzyme, namely, catalysts that can accelerate a process of interest in the thermodynamically favourable direction.
@@
Figure 1: (a) The free energy landscape that drives the internal (fuel-to-waste) thermodynamically favourable reaction of the enzyme, where $E_{\mathrm{ba}}$ is the energy barrier height and $E_{*}$ the difference in the Gibbs free energy after a complete reaction. (b) The dynamics of the preferred length of the enzyme during the internal reaction, which induces conformational changes of amplitude $\ell_{\textrm{e}}$ during a reaction. (c, d) Free energy landscapes for the molecule, representing the thermodynamically unfavourable substrate-to-product reaction, for two examples of reaction: (c) dissociation of a dimer molecule into two monomers, (d) transition of a molecule between a short and a long conformational state. Here, $\Delta$ is the barrier height, $\varepsilon$ is the Gibbs free energy difference, and $\ell_{\textrm{m}}$ is the difference in the length of molecule between the initial and the transition state.
@@
We consider an enzyme that undergoes conformational changes during the catalytic conversion of a fuel molecule to a waste molecule (Fig. 1(a,b) and Supplemental Videos 1 and 2). During each reaction, the enzyme undergoes an expansion and contraction cycle. We consider two types of substrate molecules: a dimer that can dissociate into two monomers (Fig. 1(c)), and a dimer that has two states, a short and a long one (Fig. 1(d) and Supplemental Video 3). In both cases, the dimer is attached at the outside of the enzyme. Through a detailed, thermodynamically consistent calculation, we show that during the fuel-induced contraction the enzyme can cause the dissociation of the dimer into monomers, or its transition from the short to the long conformation, respectively, even if these processes are not favoured thermodynamically or face a substantial energy barrier (see Supplemental Video 4). These catalyzed reactions are favourable in a large portion of the parameter space, related to the sizes and lengths of the enzyme and the bound molecule.
@@
Our proposed dynamical paradigm, built on appropriate implementation of the relevant physical constraints on the minimal reaction coordinates, allows us to identify the following three golden rules for the optimal function of a fuelled enzyme driven by mechanochemical coupling: (i) the enzyme and the molecule should be attached at the smaller end of each; (ii) the conformational change of the enzyme must be comparable to or larger than the conformational change required of the molecule; (iii) the conformational change of the enzyme must be fast enough so that the molecule actually stretches, rather than just following the enzyme without stretching. The rules can provide useful input to the complementary perspectives of de novo enzyme design based on machine learning and all-atom simulations.
@@ II.1 Geometry and deterministic dynamics
We start with the simple model for an enzyme that undergoes conformational changes during catalysis, as previously introduced in Ref. 35. In particular, the fuel-to-waste reaction is not modelled explicitly, and is instead described by an internal phase $\phi_{\textrm{e}}$ which moves stochastically along a downhill washboard potential $V_{\textrm{e}}(\phi_{\textrm{e}})$; see Fig. 1(a). A noise-activated jump over one of the energy barriers (of height $E_{\mathrm{ba}}$) leads to an increase of the internal phase $\phi_{\textrm{e}}$ by $2\pi$, and corresponds to the conversion of fuel into waste with a free energy release given by $E_{*}$. We assume that the catalytic process is reaction-limited, which is equivalent to fuel molecules being sufficiently abundant, such that a new fuel molecule would effectively bind instantly to the enzyme after a reaction allowing the process to be repeated again.
@@
The enzyme is represented by two sub-units, with the separation $L_{\textrm{e}}$ between them representing the mechanical degree of freedom that is coupled to the internal chemistry. This coupling is described by the potential $U(L_{\textrm{e}},\phi_{\textrm{e}})=\frac{k}{2}(L_{\textrm{e}}-L(\phi_{\textrm{e}}))^{2}+V_{\textrm{e}}(\phi_{\textrm{e}})$, where the first term is a harmonic potential that couples the actual length $L_{\textrm{e}}$ to a preferred length $L(\phi_{\textrm{e}})$ that depends on the internal phase describing the fuel-to-waste reaction; see Fig. 1(b). As a consequence of this coupling, every time that a fuel-to-waste reaction occurs the enzyme undergoes a cyclic conformational change.
@@
The dynamics of the dimer molecule is described by a single degree of freedom, which is the separation $L_{\textrm{m}}$ between its two monomers; see Fig. 1(c,d). To model the two different chemical reactions already described (dimer dissociation and short-to-long conformation switch), two different potentials $V_{\textrm{m}}(L_{\textrm{m}})$ that govern the length of the molecule are used. For convenience and to unify the notation in both cases, we express the dimer length as $L_{\textrm{m}}=L_{\textrm{m}}^{(0)}+\ell_{\textrm{m}}\phi_{\textrm{m}}$ where $L_{\textrm{m}}^{(0)}$ is the length in the dimer state or in the transition state between short and long, respectively, and $\ell_{\textrm{m}}$ is the length increase necessary to reach the transition state; see Fig. 1(c,d).
@@
Figure 2: (a) A visualization of a dissociation process catalyzed by a fuelled enzyme, divided into four stages. The blue and red curves correspond to the energy landscape of the internal reaction of the enzyme $V_{\textrm{e}}(\phi_{\textrm{e}})$, and of the molecular reaction $V_{\textrm{m}}(\phi_{\textrm{m}})$, respectively. (b) The same process visualized within the two dimensional energy landscape given by $V_{\textrm{e}}(\phi_{\textrm{e}})+V_{\textrm{m}}(\phi_{\textrm{m}})$, where the black line shows the evolution of the combined enzyme-molecule system. Thermal noise tends to kick the system over the enzymatic energy barrier, since it is smaller than the barrier in the energy landscape of the molecule. The coupling between the enzyme and molecule dynamics, mediated by the conformational changes of the enzyme, then causes the system to cross over the energy barrier of the molecular reaction.
@@
We now briefly describe the dynamics governing these degrees of freedom (for the detailed derivation see Appendix A). When the substrate molecule strongly binds to the enzyme, a complex effectively made up of three sub-units is formed (Fig. 2). Each sub-unit has a hydrodynamic mobility (inverse friction): $\mu_{\textrm{e}}$ for the part of the enzyme not bound to the molecule, $\mu_{\textrm{b}}$ for the unit that is shared by the molecule and the enzyme when bound to each other, and $\mu_{\textrm{m}}$ for the part of the molecule that is not bound to the enzyme. Since the typical time scales of the catalytic activity and the associated enzyme conformational transitions are much longer than the relevant inertial time scale [32], we consider the overdamped dynamics of these three sub-units, which leads to coupled evolution equations for the two lengths. Similarly, we consider overdamped dynamics for the evolution of the internal phase of the enzyme along the chemical free energy landscape, with an associated mobility $\mu_{\phi}$. By assuming that the enzyme is stiff, one can enslave the dynamics of the enzyme length (fast variable) onto the dynamics of the internal phase (slow variable). This reduces the degrees of freedom to two, and results in coupled equations for the internal phase of the enzyme and the length of the molecule. The deterministic components of the equations take on the form
@@ II.2 The effect of noise in the stochastic dynamics
Since enzymes operate at the nano-scale, they are strongly affected by the thermal fluctuations of the surrounding medium. In particular, because the dynamics is dissipative, thermal noise is essential to kick the system out of local energy minima and over barriers associated with the reactions; see Fig. 1(a,c,d).
Therefore, it is essential to introduce thermal fluctuations in our description of the problem. The system is out of equilibrium due to the driving force given by the energy $E_{*}$ of the fuel. In the limit of $E_{*}=0$, the steady state of the system must be a state of thermodynamic equilibrium, corresponding to a Boltzmann distribution. This implies that the stochastic dynamics corresponding to the deterministic dynamics in Eq. (1) must be described by a Fokker-Planck equation describing the time evolution of the probability distribution $P(\phi_{\textrm{e}},\phi_{\textrm{m}},t)$ as follows
@@
The Langevin equation associated to the Fokker-Planck equation in Eq. (5) takes the following form in the Stratonovich convention,
@@
where $\sum_{\nu=\textrm{e},\textrm{m}}\sigma_{\alpha\nu}\sigma_{\beta\nu}=M_{\alpha\beta}$. The term on the second line of the right hand side is the spurious drift term arising from the multiplicative nature of the noise, due to the dependence of the mobility matrix on $\phi_{\textrm{e}}$, while the last term on the first line of Eq. (7) is the thermal noise with $\xi_{\beta}$ being a unit white noise such that $\langle\xi_{\beta}(t)\rangle=0$ and $\langle\xi_{\alpha}(t)\xi_{\beta}(t^{\prime})\rangle=\delta_{\alpha\beta}\delta(t-t^{\prime})$. It is also worth noting that the mobility matrix must be symmetric and positive definite (which is the case here) in order for the dynamics to be thermodynamically consistent [50, 51].
@@ II.3 Model parametrization
The enzyme and the molecule are embedded in a solvent under low Reynolds number conditions [51]. Therefore, we can infer that the mobilities $\mu_{\textrm{e}}$ (enzyme-only sub-unit), $\mu_{\textrm{b}}$ (bound enzyme-molecule sub-unit), and $\mu_{\textrm{m}}$ (molecule-only sub-unit) are related to their corresponding effective hydrodynamic radii $a_{\textrm{e}}$, $a_{\textrm{b}}$, and $a_{\textrm{m}}$, via the Stokes relation $\mu=1/(6\pi\eta a)$, ignoring hydrodynamic interactions in this lowest order approximation. The mobilities thus directly encode information about the sizes of the enzyme and the molecule sub-units. The length scale related to the amplitude of the deformation of the molecule is $\ell_{\textrm{m}}$; see Fig. 1(c,d). For the conformational changes of the enzyme, we assume that the preferred length $L(\phi_{\textrm{e}})$ oscillates during the enzymatic cycle as
@@
The washboard potential governing the dynamics of the internal phase $\phi_{\textrm{e}}$ is taken to be of the form $V_{\textrm{e}}(\phi_{\textrm{e}})=-F\phi_{\textrm{e}}+v\sin(\phi_{\textrm{e}}-\delta)$ where $\delta=\arccos(F/v)$ is the value of the phase shift which guarantees that the minima of the potential are located at integer multiples of $2\pi$. The parameters $F$ and $v$ can be mapped to the height of the energy barrier of the enzymatic reaction $E_{\mathrm{ba}}$ and the free energy released after a complete reaction $E_{*}$ (see Fig. 1(a)) through $E_{\mathrm{ba}}=[2\sqrt{1-(F/v)^{2}}+2\delta(F/v)]v$ and $E_{*}=2\pi F$ [35, 47, 52]. By eliminating $v$ and $F$, we find that the parameter $\delta$ is given by the solution of the transcendental equation
@@
The parameter $\mu_{\phi}$ is the mobility of the internal phase $\phi_{\textrm{e}}$ and may have different microscopic origins depending on the particulars of the fuel-to-waste reaction. In principle, it can be related to the rate $k_{\mathrm{cat}}$ of the catalytic reaction and the height of the energy barrier $E_{\mathrm{ba}}$ through the Kramers escape rate [8, 9]
$k_{\mathrm{cat}}=\frac{\mu_{\phi}}{2\pi}\sqrt{\lvert\lambda\rvert\lambda^{(0)}}\exp\bigg{(}-\frac{E_{\mathrm{ba}}}{k_{\rm B}T}\bigg{)},$ (11)
where $\lambda^{(0)}=V_{\textrm{e}}^{\prime\prime}|_{\rm min}$ and $\lambda=V_{\textrm{e}}^{\prime\prime}|_{\rm max}$, i.e. they correspond to the values of the second derivative (curvature) of the potential around the minimum and the maximum (energy barrier), respectively. Thus, by knowing the catalytic rate and barrier height of a reaction one can infer the value of $\mu_{\phi}$.
For the molecule, we consider two different types of reaction starting from a dimer-like substrate. For the dissociation reaction, we consider a potential with a minimum representing the stable initial state of the molecule, a maximum corresponding to an energy barrier with height $\Delta$, and an asymptotic decrease to a constant value, which represents the fact that beyond the barrier the monomers become disconnected. We describe this with the potential $V_{\textrm{m}}(\phi_{\textrm{m}})=\Delta\phi_{\textrm{m}}^{2}\exp[2(1-\phi_{\textrm{m}})]$, shown in Fig. 1(c). For the conformational switch reaction where the dimer switches from a short to a long state, we choose a bistable potential, with an energy barrier $\Delta$ and an energy difference $\varepsilon$ between the long (higher energy) and the short (lower energy) states. We model this using the potential $V_{\textrm{m}}(\phi_{\textrm{m}})=\Delta\phi_{\textrm{m}}^{2}(\phi_{\textrm{m}}^{2}-2)+\varepsilon\phi_{\textrm{m}}/2$, with $\varepsilon<\Delta$; see Fig. 1(d). All the relevant model parameters are summarized in Table 1.
Figure 3: Phase-portraits of the deterministic dynamics for (a,b) the dissociation reaction and (c,d) the switch reaction. A global bifurcation occurs with increasing coupling strength $h$, which is below the critical value in (a,c) and above it in (b,d). In all cases, stable fixed points are depicted as circles, saddle points as diamonds, and unstable fixed points as asterisks. Dashed lines depict possible scenarios in which thermal fluctuations can kick the system out of the basin of attraction of a stable fixed point. The solid flow lines show typical trajectories, and their final resting points, when starting from different basins of attraction. The different colouring of the diamonds is described in the text. The values of the coupling are (a) $h=0.33$, (b) $h=0.48$, (c) $h=0.12$, and (d) $h=0.48$.

### v2

# Mechanistic rules for de novo design of enzymes

@@ Abstract
Enzymes are nano-scale machines that have evolved to drive chemical reactions out of equilibrium in the right place at the right time. Given the complexity and specificity of enzymatic function, bottom-up design of enzymes presents a daunting task that is far more challenging than making passive molecules with specific binding affinities or building nano-scale mechanically active devices. We present a thermodynamically-consistent model for the operation of such a fuelled enzyme, which uses the energy from a favourable reaction to undergo non-equilibrium conformational changes that in turn catalyze a chemical reaction on an attached substrate molecule. We show that enzymatic function can emerge through a bifurcation upon appropriate implementation of momentum conservation on the effective reaction coordinates of the low dimensional description of the enzyme, and thanks to a generically present dissipative coupling. Our results can complement the recently developed strategies for de novo enzyme design based on machine learning approaches.
@@ I Introduction
In recent years, fundamental insight from statistical physics has been successfully used towards experimental demonstrations of many interesting non-equilibrium phenomena [15, 16, 17, 18]. To aim towards the development of systems that can accomplish relatively complex tasks, one can benefit from innovative ideas that build on emergent physical properties of many-body non-equilibrium systems. Recent examples of such developments include proposals to employ non-reciprocal interactions to design collective barrier-crossing strategies that could emulate enzymatic function at larger scales [19, 20], autonomous multifarious self-organization of complex protein structures [21], as well as proposed scenarios for fast and efficient self-organization of primitive metabolic cycles at the early stages of life formation [22].
@@
Here, we set out to construct a minimal model for a ‘fuelled’ enzyme, so that we can extract a set of golden rules regarding its optimal design (see below). Our theoretical framework is built upon two main pillars, namely, momentum conservation as a fundamental physical constraint on the effective reaction coordinates that span the relevant low dimensional configuration space of the system, and a generically present dissipative coupling between the different reaction coordinates, which we systematically derive from a microscopic model. The implementation of momentum conservation leads to the emergence of an effective mechanochemical coupling in the same spirit as models used in studies of stochastic nano-swimmers [30, 31, 32, 33] and enhanced diffusion of enzymes [34, 35], while the dissipative coupling has been recently used to show how enzyme pairs can cooperate by exhibiting synchronization and enhanced catalytic activity [36]. The two ingredients described above give rise to a bifurcation in the dynamical phase-portrait of the low dimensional configuration space, which enables enzymatic activity as an emergent feature of the dynamics, via a fundamentally novel mechanism not accessible to Kramers-like energy-barrier-crossing descriptions in terms of reaction coordinates. Typically, those descriptions are based on a modification of the energy landscape of the reaction induced by the presence of the enzyme. Here, the energy landscape remains unchanged in the presence of the enzyme, but the dynamics on this landscape becomes qualitatively different due to the cross-talk between different reaction coordinates which arises naturally as a result of the bifurcation. Our work complements other studies that involve minimal models for studying enzymatic behaviour, often based on colloids interacting through short-ranged interaction potentials or bead-spring networks [37, 38, 39, 40, 41, 42, 43]. These studies predominantly describe the action of a ‘passive’ enzyme, namely, catalysts that can accelerate a process of interest in the thermodynamically favourable direction.
@@
Figure 1: Minimal model of an enzymatic reaction (A) We focus only on the catalytic step of the reaction, where a fuel-to-waste conversion ($\mathrm{\phi_{\textrm{e}}}$) causes a substrate-to-product reaction. For simplicity, we omit the substrate binding ($k_{\mathrm{on}}$) and product unbinding ($k_{\mathrm{off}}$) steps, assuming that they take place rapidly after each catalytic reaction. (B) The free energy landscape that drives the internal (fuel-to-waste) thermodynamically favourable reaction of the enzyme, where $E_{\mathrm{ba}}$ is the energy barrier height and $E_{*}$ the difference in the Gibbs free energy after a complete reaction. (C) The dynamics of the preferred length of the enzyme during the internal reaction, which induces conformational changes of amplitude $\ell_{\textrm{e}}$ during a reaction. (D, E) Free energy landscapes for the molecule, representing the thermodynamically unfavourable substrate-to-product reaction, for two examples of reaction: (D) dissociation of a dimer molecule into two monomers, (E) transition of a molecule between a short and a long conformational state. Here, $\Delta$ is the barrier height, $\varepsilon$ is the Gibbs free energy difference, and $\ell_{\textrm{m}}$ is the difference in the length of molecule between the initial and the transition state.
@@
We focus on the catalytic step of an enzymatic reaction (Figure 1(A)) and consider the conformational changes of the enzyme during the conversion of a fuel molecule to a waste molecule (Figure 1(B,C) and Supplemental Videos 1 and 2). During each reaction, the enzyme undergoes an expansion and contraction cycle. This could also serve as a simple model of an active nanomotor that undergoes shape oscillations, as used in recent experiments [44]. We consider two types of substrate molecules: a dimer that can dissociate into two monomers (Figure 1(D)), and a dimer that has two states, a short and a long one (Figure 1(E) and Supplemental Video 3). In both cases, the dimer is attached at the outside of the enzyme. Through a detailed, thermodynamically consistent calculation, we show that during the fuel-induced contraction the enzyme can cause the dissociation of the dimer into monomers, or its transition from the short to the long conformation, respectively, even if these processes are not favoured thermodynamically or face a substantial energy barrier (see Supplemental Video 4). These catalyzed reactions are favourable in a large portion of the parameter space, related to the sizes and lengths of the enzyme and the bound molecule.
@@
Our proposed dynamical paradigm, built on appropriate implementation of the relevant physical constraints on the minimal reaction coordinates, allows us to identify the following three golden rules for the optimal function of a fuelled enzyme driven by mechanochemical coupling: (i) the enzyme and the molecule should be attached at their corresponding smaller ends (i.e. higher mobility ends); (ii) the conformational change of the enzyme must be comparable to, or larger than, the conformational change required of the molecule; (iii) the conformational change of the enzyme must be sufficiently fast such that the molecule is stretched, rather than simply translated along with the enzyme without stretching. The rules can provide useful input to the complementary perspectives of de novo enzyme design based on machine learning and all-atom simulations.
@@ II.1 Geometry and deterministic dynamics
We start with the simple model for an enzyme that undergoes conformational changes during a fuel-to-waste reaction, as previously introduced in Ref. 36. In particular, the fuel-to-waste reaction is not modelled explicitly, and is instead described by an internal phase $\phi_{\textrm{e}}$ which moves stochastically along a downhill washboard potential $V_{\textrm{e}}(\phi_{\textrm{e}})$; see Figure 1(B). A noise-activated jump over one of the energy barriers (of height $E_{\mathrm{ba}}$) leads to an increase of the internal phase $\phi_{\textrm{e}}$ by $2\pi$, and corresponds to the conversion of fuel into waste with a free energy release given by $E_{*}$. We assume that the fuel-to-waste process is reaction-limited, which is equivalent to fuel molecules being sufficiently abundant, such that a new fuel molecule would effectively bind instantly to the enzyme after a reaction allowing the process to be repeated again (see Figure 1(A)).
@@
The enzyme is represented by two sub-units, with the separation $L_{\textrm{e}}$ between them representing the mechanical degree of freedom that is coupled to the internal chemistry. This coupling is described by the potential $U(L_{\textrm{e}},\phi_{\textrm{e}})=\frac{k}{2}(L_{\textrm{e}}-L(\phi_{\textrm{e}}))^{2}+V_{\textrm{e}}(\phi_{\textrm{e}})$, where the first term is a harmonic potential that couples the actual length $L_{\textrm{e}}$ to a preferred length $L(\phi_{\textrm{e}})$ that depends on the internal phase describing the fuel-to-waste reaction; see Figure 1(C). As a consequence of this coupling, every time that a fuel-to-waste reaction occurs the enzyme undergoes a cyclic conformational change.
@@
The dynamics of the dimer molecule is described by a single degree of freedom, which is the separation $L_{\textrm{m}}$ between its two monomers; see Figure 1(D,E). To model the two different chemical reactions already described (dimer dissociation and short-to-long conformation switch), two different potentials $V_{\textrm{m}}(L_{\textrm{m}})$ that govern the length of the molecule are used. For convenience and to unify the notation in both cases, we express the dimer length as $L_{\textrm{m}}=L_{\textrm{m}}^{(0)}+\ell_{\textrm{m}}\phi_{\textrm{m}}$ where $L_{\textrm{m}}^{(0)}$ is the length in the dimer state or in the transition state between short and long, respectively, and $\ell_{\textrm{m}}$ is the length increase necessary to reach the transition state; see Figure 1(D,E).
@@
Figure 2: The evolution along the two-dimensional energy landscape (A) A visualization of a dissociation process catalyzed by a fuelled enzyme, divided into four stages. The blue and red curves correspond to the energy landscape of the internal reaction of the enzyme $V_{\textrm{e}}(\phi_{\textrm{e}})$, and of the molecular reaction $V_{\textrm{m}}(\phi_{\textrm{m}})$, respectively. The solid black circles on the energy landscapes indicate the present state of each process, depicted in each respective schematic figure. The red arrows represent internal forces on the enzyme and molecule. (B) The same process visualized within the two dimensional energy landscape given by $V_{\textrm{e}}(\phi_{\textrm{e}})+V_{\textrm{m}}(\phi_{\textrm{m}})$, where the black line shows the evolution of the combined enzyme-molecule system. Thermal noise tends to kick the system over the enzymatic energy barrier, since it is smaller than the barrier in the energy landscape of the molecule. The coupling between the enzyme and molecule dynamics, mediated by the conformational changes of the enzyme, then causes the system to cross over the energy barrier of the molecular reaction.
@@
We now briefly describe the dynamics governing these degrees of freedom (for the detailed derivation see SI). When the substrate molecule strongly binds to the enzyme, a complex effectively made up of three sub-units is formed (Figure 2). Each sub-unit has a hydrodynamic mobility (inverse friction): $\mu_{\textrm{e}}$ for the part of the enzyme not bound to the molecule, $\mu_{\textrm{b}}$ for the unit that is shared by the molecule and the enzyme when bound to each other, and $\mu_{\textrm{m}}$ for the part of the molecule that is not bound to the enzyme. Since the typical time scales of the catalytic activity and the associated enzyme conformational transitions are much longer than the relevant inertial time scale [33], we consider the overdamped dynamics of these three sub-units, which leads to coupled evolution equations for the two lengths. Similarly, we consider overdamped dynamics for the evolution of the internal phase of the enzyme along the chemical free energy landscape, with an associated mobility $\mu_{\phi}$. By assuming that the enzyme is stiff, one can enslave the dynamics of the enzyme length (fast variable) onto the dynamics of the internal phase (slow variable). This reduces the degrees of freedom to two, and results in coupled equations for the internal phase of the enzyme and the length of the molecule. The deterministic components of the equations take on the form
@@ II.2 The effect of noise in the stochastic dynamics
Since enzymes operate at the nano-scale, they are strongly affected by the thermal fluctuations of the surrounding medium. In particular, because the dynamics is dissipative, thermal noise is essential to kick the system out of local energy minima and over barriers associated with the reactions; see Figure 1(B,D,E).
Therefore, it is essential to introduce thermal fluctuations in our description of the problem. The system is out of equilibrium due to the driving force given by the energy $E_{*}$ of the fuel. In the limit of $E_{*}=0$, the steady state of the system must be a state of thermodynamic equilibrium, corresponding to a Boltzmann distribution. This implies that the stochastic dynamics corresponding to the deterministic dynamics in Equation (1) must be described by a Fokker-Planck equation describing the time evolution of the probability distribution $P(\phi_{\textrm{e}},\phi_{\textrm{m}},t)$ as follows
@@
The Langevin equation associated to the Fokker-Planck equation in Equation (5) takes the following form in the Stratonovich convention,
@@
where $\sum_{\nu=\textrm{e},\textrm{m}}\sigma_{\alpha\nu}\sigma_{\beta\nu}=M_{\alpha\beta}$. The term on the second line of the right hand side is the spurious drift term arising from the multiplicative nature of the noise, due to the dependence of the mobility matrix on $\phi_{\textrm{e}}$, while the last term on the first line of Equation (7) is the thermal noise with $\xi_{\beta}$ being a unit white noise such that $\langle\xi_{\beta}(t)\rangle=0$ and $\langle\xi_{\alpha}(t)\xi_{\beta}(t^{\prime})\rangle=\delta_{\alpha\beta}\delta(t-t^{\prime})$. It is also worth noting that the mobility matrix must be symmetric and positive definite (which is the case here) in order for the dynamics to be thermodynamically consistent [53, 54].
@@ II.3 Model parametrization
The enzyme and the molecule are embedded in a solvent under low Reynolds number conditions [54]. Therefore, we can infer that the mobilities $\mu_{\textrm{e}}$ (enzyme-only sub-unit), $\mu_{\textrm{b}}$ (bound enzyme-molecule sub-unit), and $\mu_{\textrm{m}}$ (molecule-only sub-unit) are related to their corresponding effective hydrodynamic radii $a_{\textrm{e}}$, $a_{\textrm{b}}$, and $a_{\textrm{m}}$, via the Stokes relation $\mu=1/(6\pi\eta a)$, ignoring hydrodynamic interactions in this lowest order approximation. The mobilities thus directly encode information about the sizes of the enzyme and the molecule sub-units. The length scale related to the amplitude of the deformation of the molecule is $\ell_{\textrm{m}}$; see Figure 1(D,E). For the conformational changes of the enzyme, we assume that the preferred length $L(\phi_{\textrm{e}})$ oscillates during the enzymatic cycle as
@@
The washboard potential governing the dynamics of the internal phase $\phi_{\textrm{e}}$ is taken to be of the form $V_{\textrm{e}}(\phi_{\textrm{e}})=-F\phi_{\textrm{e}}+v\sin(\phi_{\textrm{e}}-\delta)$ where $\delta=\arccos(F/v)$ is the value of the phase shift which guarantees that the minima of the potential are located at integer multiples of $2\pi$. The parameters $F$ and $v$ can be mapped to the height of the energy barrier of the enzymatic reaction $E_{\mathrm{ba}}$ and the free energy released after a complete reaction $E_{*}$ (see Figure 1(B)) through $E_{\mathrm{ba}}=[2\sqrt{1-(F/v)^{2}}+2\delta(F/v)]v$ and $E_{*}=2\pi F$ [36, 50, 55]. By eliminating $v$ and $F$, we find that the parameter $\delta$ is given by the solution of the transcendental equation
@@
The parameter $\mu_{\phi}$ is the mobility of the internal phase $\phi_{\textrm{e}}$ and may have different microscopic origins depending on the particulars of the fuel-to-waste reaction. In principle, it can be related to the rate $k(\mathrm{f}\to\mathrm{w})$ of the fuel-to-waste reaction and the height of the energy barrier $E_{\mathrm{ba}}$ through the Kramers escape rate [8, 9]
$k(\mathrm{f}\to\mathrm{w})=\frac{\mu_{\phi}}{2\pi}\sqrt{\lvert\lambda_{\mathrm{e}}^{(1)}\rvert\lambda_{\mathrm{e}}^{(0)}}\exp\bigg{(}-\frac{E_{\mathrm{ba}}}{k_{\rm B}T}\bigg{)},$ (11)
where $\lambda_{\mathrm{e}}^{(0)}=V_{\textrm{e}}^{\prime\prime}(\phi_{\mathrm{e}})|_{\rm min}$ and $\lambda_{\mathrm{e}}^{(1)}=V_{\textrm{e}}^{\prime\prime}(\phi_{\mathrm{e}})|_{\rm max}$, i.e. they correspond to the values of the second derivative (curvature) of the enzyme potential around the minimum and the maximum (energy barrier), respectively. Thus, by knowing the reaction rate and barrier height one can infer the value of $\mu_{\phi}$.
For the molecule, we consider two different types of reaction starting from a dimer-like substrate. For the dissociation reaction, we consider a potential with a minimum representing the stable initial state of the molecule, a maximum corresponding to an energy barrier with height $\Delta$, and an asymptotic decrease to a constant value, which represents the fact that beyond the barrier the monomers become disconnected. We describe this with the potential $V_{\textrm{m}}(\phi_{\textrm{m}})=\Delta\phi_{\textrm{m}}^{2}\exp[2(1-\phi_{\textrm{m}})]$, shown in Figure 1(D). For the conformational switch reaction where the dimer switches from a short to a long state, we choose a bistable potential, with an energy barrier $\Delta$ and an energy difference $\varepsilon$ between the long (higher energy) and the short (lower energy) states. We model this using the potential $V_{\textrm{m}}(\phi_{\textrm{m}})=\Delta\phi_{\textrm{m}}^{2}(\phi_{\textrm{m}}^{2}-2)+\varepsilon\phi_{\textrm{m}}/2$, with $\varepsilon<\Delta$; see Figure 1(E). All the relevant model parameters are summarized in Table 1.
Figure 3: Phase-portraits of the deterministic dynamics (A,B) the dissociation reaction and (C,D) the switch reaction. A global bifurcation occurs with increasing coupling strength $h$, which is below the critical value in (A,C) and above it in (B,D). In all cases, stable fixed points are depicted as circles, saddle points as diamonds, and unstable fixed points as asterisks. Dashed lines depict possible scenarios in which thermal fluctuations can kick the system out of the basin of attraction of a stable fixed point. The solid flow lines show typical trajectories, and their final resting points, when starting from different basins of attraction. The different colouring of the diamonds is described in the text. The values of the coupling are (A) $h=0.33$, (B) $h=0.48$, (C) $h=0.12$, and (D) $h=0.48$.